OpenStudy (anonymous):

i got problem with differential. i know that (e^x)'=e^x but why (e^4x)'=4e^x ?? whyn not 4e^x-1 like all the other equations? its just an exception?

OpenStudy (anonymous):

\[\frac{d}{dx}e^{f(x)}=f'(x)\cdot e^{f(x)}\] In this case, \(f(x)=4x\) and so \(f'(x)=4\), then \[\frac{d}{dx}e^{4x}=4e^{4x},~\text{not }4e^x\]

OpenStudy (unklerhaukus):

It's a case of the derivative of a composition \[(u\circ v)'= (u'\circ v)\cdot v'\\ %(u(v))'=u'(v)\cdot v'\]\[(e^{f})'\\ %=\exp(f(x))'\\ \qquad\qquad\qquad\qquad u(f)=e^f\qquad v=f\\ \qquad\qquad\qquad\qquad u'(f)=e^f\qquad v'=f'\\ (e^{f})'=e^{f}\cdot f'\]